OpenStudy (anonymous):

Write the equation of the ellipse in standard form9x^2+4y^2-72x+40y+208

OpenStudy (tkhunny):

1) That's not an ellipse. It's just an algebraic expression. 2) If you mean "= 0", then we can talk. 9(x^2 - 8x + _____) + 4(y^2 + 10y + _____) = -208 + 9(_____) + 4(_____)

OpenStudy (anonymous):

9x^2+4y^2-72x+40y+208=0

OpenStudy (tkhunny):

:-) Notation means things. it enables communcation if nothing else. Complete the square on x and y and let's see what you get.